# Fast-Poly: A Fast Polyhedral Framework For 3D Multi-Object Tracking

### v1

@@ I INTRODUCTION
  • Object is rotated in 3D space. Geometry-based metrics (IoU [10, 11], GIoU [2, 1, 4]) provide high interpretability and accurately capture inter-object affinity, are widely utilized. However, according to statistics in Table II, 3D object rotation leads to time-consuming polygon intersection and convex hull solution steps in these metrics. Some works [3, 12, 13] employ custom Euclidean distance to reduce overhead, but accuracy is lost due to the lack of spatial relationship considerations.
  • The tracking pipeline is from a global perspective. Modeling global object similarity for pre-processing and association is a common practice in the field [12, 2, 4, 10, 13, 11, 1, 3]. However, inter-observation redundancy and trajectory-observation correlation are only feasible in a limited local space, leading to numerous invalid computations. [1, 2, 4] emphasize restricting false positive (FP) agents from influencing cost computations. [1] also applies multiple criteria to avoid distinct categories matching. Despite these efforts, the fact that distant objects are unrelated is widely ignored.
@@
  • Matrix calculations in filters are heavy. Trajectories typically represented as high-dimensional vectors [12, 2, 4, 10, 13, 11, 1, 3]. Nevertheless, integrating time-invariant states in the Kalman Filter (KF) leads to redundant computations involving zeros. [14] accelerates 2D position estimation through a priori split calculation and CUDA operator. However, in 3D MOT, the abundance of states and non-linear motion models [1] make a similar matrix-to-quadratic form transformation challenging. Furthermore, rigid count-based lifecycle management hinders real-time performance due to matrix operations applied to all active tracklets, including the coasted FP tracklets.
@@ II Related Work
Current 3D MOT [15, 1, 2, 16, 17, 12, 5, 3, 4, 13, 11, 10, 6, 18, 19, 20, 7] benefits from advancements in 2D MOT [21, 14, 22, 23, 24] and 3D detection [25, 26, 27, 28, 20, 29, 30]. AB3DMOT [10] pioneers the expansion of the TBD framework into 3D space, introducing a simple yet effective baseline. Poly-MOT [1] proposes category-specific ideas to introduce a strong tracker based on multiple non-linear models and similarity metrics, achieving SOTA performance on nuScenes. With Neural Network (NN) and a post-processing manner, some methods [5, 6, 16, 7] utilize the features output by the pre-trained backbone [20] to regress the affinity or even the tracking confidence [6]. Leveraging the occlusion robustness of LiDAR and the far-sightedness of cameras, EagerMOT [3] performs a detection-level fusion to combine 2D and 3D detectors. CAMO-MOT [2] introduces an occlusion head to classify the visual occlusion status and implements a confidence-involved LiDAR association. JMOT [31] and mmMOT [32] develop a feature-level fusion multi-modal tracker, generating detections and association based on the Joint Detection and Tracking (JDT) framework. Being economical and an end-to-end pipeline, camera-only trackers [18, 19] based on Bird’s-Eye View (BEV) representation receive attention and applications. Following the Tracking-By-Attention (TBA) framework, these trackers implement spatial-temporal modeling based on self/cross-attention and implicitly learn trajectory motion in a self-refinement paradigm. The trade-off between computational overhead, real-time performance, and tracking accuracy stands as the primary research direction in the current landscape.
@@ II-B Tracking-By-Detection
Estimation. This module utilizes filters (Linear Kalman Filter [12, 3, 2, 4, 13, 11, 10], Extended Kalman Filter [1], Point Filter [15, 20], etc.), along with motion models, to perform two functions: (1) Predict the alive tracklets to achieve temporal alignment and association with the detection. (2) Update the matched tracklets with the corresponding observations, preparing prior information for downstream. Nevertheless, the real-time challenge emerges when managing numerous tracklets, as the filter involves heavy matrix operations on each tracklet. Therefore, we decouple and filter time-invariant states using a lightweight filter to alleviate the computational burden.
Association. As the core of the system, this module establishes tracklet-observation similarity and resolves matching correspondence. Geometry-based (IoU [10, 11], GIoU [2, 1, 4], Euclidean [15, 13, 12, 3], NN distance [16, 6, 7, 5], etc.) and appearance-based are the commonly used affinity metrics. The former modeling process, reliant solely on spatial information, is mostly unsupervised and robust to occlusion. The latter, robust to inaccurate depth, can describe distant objects. Under the one-to-one assumption, the optimal assignment [33] algorithm is then used to obtain matching pairs. However, rotation in 3D space impairs the parallelization of IoU-like metrics, introducing severe latency. Fast-Poly proposes to leverage alignment to solve this dilemma.
Lifecycle. This module initializes, terminates, and merges tracklets based on the count-based strategy [12, 3, 2, 1, 4, 11, 10] or the confidence-based [15, 1, 4, 13] strategy. The former has simple logic and strong robustness. The latter is more flexible in processing based on heuristic functions. In response to the inherent rigidity issue of the former approach, we employ score refinement to flexible terminate trajectories. In addition, we leverage trajectory average scores and a novel prediction to bolster occlusion robustness and enhance the estimation accuracy of the confidence-based strategy.
Figure 2: The pipeline of our proposed method Fast-Poly. Our structure design is illustrated in Section III-A. Real-time improvements to the baseline [1] are highlighted in distinct colors. Orange denotes the alignment to counter object rotation. Blue denotes the densification to increase computational efficiency. Cyan denotes the parallelization to alleviating serial bottleneck.
@@ III-A Overall Architecture
Despite sharing a similar pipeline, Fig. 2 illustrates the distinctions between our proposed method and Poly-MOT [1]. At each frame $t$, Fast-Poly employs two independent computing processes (Section III-D) to filter the 3D detections $D_{t}$ and predict existing trajectories $T_{t,t-1}$. Specifically, SF and NMS filters are leveraged to process the detections. Besides, the motion (time-variable), score (Fig. 3), and time-invariant (Fig. 3) states of trajectories are predicted by corresponding filters. $D_{t}$ and $T_{t,t-1}$ are then subjected to construct cost matrices $C_{t}^{1},C_{t}^{2}$ for two-stage association. Notably, during NMS and matching, our proposed voxel mask (Fig. 3) and a novel geometry-based metric (Section III-B) accelerate computing. Hungarian algorithm [33] is then employed to obtain matched pairs $DT_{t}$, unmatched detections $D_{t}^{um}$, unmatched tracklets $T_{t-1}^{um}$. To relieve the matrix dimension, time-invariant and time-variable states in matched tracklets $T_{t-1}^{m}$ are updated with corresponding observations $D_{t}^{m}$ based on our proposed lightweight filter (Fig. 3) and Extended Kalman Filter (EKF), respectively. $T_{t-1}^{m}$ scores are also refined leveraging the confidence-count mixed lifecycle (Fig. 3). $D_{t}^{um}$ are then initialized as newly active tracklets. To flexibly determine FP existing agents in $T_{t-1}^{um}$, we soft-terminate the mismatch tracklets by considering both their max-age and their online average refined scores. The remaining tracklets $T_{t}$ are sent to downstream tasks and implemented for next frame tracking.
@@ III-B Alignment
In this section, we provide a comprehensive explanation of how Fast-Poly effectively addresses high latency caused by object rotations through alignment, covering both implementation and underlying principles.
@@
A-GIoU. Varied from the pixel plane, the rotation causes overlap and convex hull irregular polygons instead of axis-aligned rectangles. The complexity of the well-known Sutherland-Hodgman [34] and Graham scan [35] algorithm are $\mathcal{O}(n^{2})$ and $\mathcal{O}(n\log{n})$, where $n$ is the number of input points. In addition, the sum calculation time increases rapidly with the growth of detections and tracklets. On the contrary, the solution complexity of the axis-aligned bounding box is only $\mathcal{O}(1)$. To this end, we proposed the Aligned Generalized Intersection over Union ($A\text{-}gIoU$), which is formulated as:
@@
where $B^{A}$ represents the axis-aligned $B$, specially containing the top-left and bottom-right points $p\in\mathbb{R}^{1\times 4}$ on the ground plane. Using the BEV space as a case study, the alignment procedure is depicted in Fig. 3. The overlap and convex hull calculations in Eq. 2 are consistent with $gIoU_{2d}$ [36], with $\mathcal{O}(1)$ computational complexity.
Naturally, two critical questions arise: (1) Can $A\text{-}gIoU$ effectively capture inter-object similarity? Our answer is affirmative. As demonstrated in Tables VI and II, on nuScenes, integrating $A\text{-}gIoU$ in association reduces latency by 73%, while even enhancing accuracy (+0.5% MOTA). Furthermore, (2) why does it perform well in the tracking task? From a single tracklet or detection perspective, due to the symmetry [36] of $gIoU$, the orientation error between $B^{A}$ and $B$ is $\leq 45^{\circ}$. The limited angular loss indicates that $B^{A}$ can still characterize the spatial information of the object. From the perspective of the correlation between detections and tracklets, Fig. 3 demonstrates that the distances of matched pairs remain unchanged (or changed slightly), thereby illustrating how $A\text{-}gIoU$ works. Intuitively, the alignment introduces a uniform scale offset on each trajectory and its corresponding observation. Leveraging the size invariance [36] of $gIoU$, this consistent scale shift preserves the final similarity. Consequently, the matched similarity remains notably higher than that of false matches. Such consistency guarantees the Hungarian algorithm produces similar assignment structures compared to $gIoU$.
@@ III-C Densification
In this section, we improve the computational efficiency of the 3D tracking system through three key enhancements: softly eliminating the coasted FP tracklets, introducing the voxel mask, and splitting time-invariant states.
Figure 3: Top: the calculation process of distinct metrics in the BEV space. Bottom: the similarity of matched pair between the identical trajectory and detection across two random frames using original $gIoU$ and our proposed $A\text{-}gIoU$.
Confidence-Count Mixed Lifecycle. The classic count-based lifecycle seriously raises tracker maintenance costs. In contrast, [15] flexibly terminates tracklets by manipulating scores, significantly reducing FP. However, as evidenced in [15] and Table VII, the deletion based on the latest score and the application of minus prediction detrimentally affect mismatch tracklets due to occlusion or FN detections. Towards this, a power function is first implemented to predict the scores smoothly. The refinement is described in two steps as follows:
@@
where $s_{t-1}$, $s_{t,t-1}$ and $s_{t}$ are the $t-1$ frame posterior score, $t$ frame priori score and $t$ frame posterior score of current tracklet, respectively. $c_{t}$ is the corresponding detection score. $\sigma$ is the hand-crafted decay rate, empirically $\leq 0.7$. The update function is similar with [15], satisfies that $s_{t}$ is more confidence than $s_{t,t-1}$ or $c_{t}$. After refinement, tracklets are terminated if their online average scores (computed from the initial frame to the current frame) fall below the deletion threshold $\theta_{dl}$ or if the number of mismatch frames exceeds the max-age. There are two noteworthy benefits: (1) Tracklets demonstrate increased resilience to temporary occlusions since their average confidence score decays gradually, as depicted in Table VII. (2) The power function swiftly pulls the tracklet predict score to the $\left[0,\sigma\right]$ interval, leaving the over-confidence area ($\left[0.7,1\right]$ interval in Fig. 2 of [15]). This reduces the hysteresis in score estimates, improving the tracking performance.
Voxel Mask. In 3D MOT, an overlooked consideration is: redundancy and matching are improbable between objects separated by significant distances in 3D Euclidean space. This stands in stark contrast to inherently no-depth 2D MOT [14, 21, 22]. To address this issue, we propose the voxel mask inspired by the coarse-to-fine idea [37, 38]. Specifically, before computing the global cost based on Eq. 2, i.e., data association and NMS iteration, we employ low-overhead Euclidean distance (2-norm) $||\cdot||_{2}$ and hand-crafted size $\theta_{vm}$ to generate voxel mask. The process can be expressed as:
@@
where $\mathcal{B}$ is the 3D bounding boxes set representing detections or tracklets based on module representation. $\mathcal{B}^{xyz}$ are the 3D center positions. $M_{t}$ is the voxel mask between $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. $\text{Aff}(\cdot)$ is the affinity function, i.e., Eq. 2 for association and $IoU_{bev}$ for pre-processing in Fast-Poly. Notably, $\text{Aff}(\cdot)$ directly pads invalid values into false index within $M_{t}$ to expedite, ultimately yielding $C_{t}$.
Lightweight Filter. Current 3D MOT methods embed the time-invariant state (z-position $z$, width $w$, length $l$, height $h$) of trajectory into the Kalman Filter. However, the time invariance makes the Jacobian matrices of process and observation identity matrices, resulting in numerous meaningless 0-related calculations. Besides, the computational time escalates swiftly with an expanding number of trajectories. To this end, we decouple these states $X^{ti}$ from the KF and estimate them using a compact filter. The estimate process is expressed as:
@@
where $X^{ti}_{t-1}$, $X^{ti}_{t,t-1}$ and $X^{ti}_{t}$ are the $t-1$ frame posterior $X^{ti}$, $t$ frame priori $X^{ti}$ and $t$ frame posterior $X^{ti}$ of current tracklet, respectively. $D^{i}_{t}$ is the corresponding detection. LW-Filter($\cdot$) is the custom lightweight filter, implemented by a median or mean filter with length $l_{lw}$. There are two key insights: (1) LW-Filter($\cdot$) efficiently circumvents invalid computations, facilitating rapid filtering while preserving temporal information. (2) In contrast to KF, which entails manual process and observation noise, LW-Filter($\cdot$) relies solely on a single hyperparameter $l_{lw}$, reducing overfitting and ensuring robustness, as proven in Fig. 5.
Input: Current 3D raw detections $D_{t}^{\prime}$; Previous alive trajectories $T_{t-1}$; Frame number $N_{f}$
Output: Alive trajectories at current frame $T_{t}$
1 Initialization: Two parallel processes $P_{1},P_{2}$; Shared flag $Flag=1$; Shared variables $V=\emptyset$ Process $P_{1}$: Pre-processing raw detections of each frame $t$ for $t$ normal-$\leftarrow$ 00 to $N_{f}$ do
2       wait $Flag==1$ then
3             $D_{t}$leftarrow Pre-processing($D_{t}^{\prime}$) $V$leftarrow $D_{t}$; $Flag$leftarrow 00
4       end
5
6 end for
7# Process $P_{2}$: Predict trajectories and implement online tracking for $t$ normal-$\leftarrow$ 00 to $N_{f}$ do
8       $T_{t,t-1}$leftarrow Predict($T_{t-1}$) wait $Flag==0$ then
9             $T_{t}$leftarrow Tracking($T_{t,t-1},V(D_{t})$) $V$leftarrow $\emptyset$; $Flag$leftarrow $1$
10       end
11
12 end for
Return: $T_{t},t=0,1,...,N_{f}$
Algorithm 1 Pseudo-code of Parallelization

### v2

@@ I INTRODUCTION
  • Object is rotated in 3D space. Geometry-based metrics (IoU [10, 11], GIoU [2, 1, 4]) provide high interpretability and accurately capture inter-object affinity, are widely utilized. However, according to statistics in Table II, 3D object rotation leads to time-consuming polygon intersection and convex hull solution steps in these metrics. Some works [3, 12] employ custom Euclidean distance to reduce overhead, but accuracy is lost due to the lack of spatial relationship considerations.
  • The tracking pipeline is from a global perspective. Modeling global object similarity for pre-processing and association is a common practice in the field [2, 4, 10, 12, 11, 1, 3]. However, inter-observation redundancy and trajectory-observation correlation are only feasible in a limited local space, leading to numerous invalid computations. [1, 2, 4] emphasize restricting false positive (FP) agents from influencing cost computations. [1] also applies multiple criteria to avoid distinct categories matching. Despite these efforts, the fact that distant objects are unrelated is widely ignored.
@@
  • Matrix calculations in filters are heavy. Trajectories typically represented as high-dimensional vectors [2, 4, 10, 12, 11, 1, 3]. Nevertheless, integrating time-invariant states in the Kalman Filter (KF) leads to redundant computations involving zeros. [13] accelerates 2D position estimation through a priori split calculation and CUDA operator. However, in 3D MOT, the abundance of states and non-linear motion models [1] make a similar matrix-to-quadratic form transformation challenging. Furthermore, rigid count-based lifecycle management hinders real-time performance due to matrix operations applied to all active tracklets, including the coasted FP tracklets.
@@ II Related Work
Current 3D MOT [14, 1, 2, 15, 16, 5, 3, 4, 12, 11, 10, 6, 17, 18, 19, 7] benefits from advancements in 2D MOT [13, 20, 21, 22] and 3D detection [23, 24, 25, 19, 26, 27]. AB3DMOT [10] pioneers the expansion of the TBD framework into 3D space, introducing a simple yet effective baseline. Poly-MOT [1] proposes category-specific ideas to introduce a strong tracker based on multiple non-linear models and similarity metrics, achieving SOTA performance on nuScenes. With Neural Network (NN) and a post-processing manner, some methods [5, 6, 15, 7] utilize the features output by the pre-trained backbone [19] to regress the affinity or even the tracking confidence [6]. Leveraging the occlusion robustness of LiDAR and the far-sightedness of cameras, EagerMOT [3] performs a detection-level fusion to combine 2D and 3D detectors. CAMO-MOT [2] introduces an occlusion head to classify the visual occlusion status and implements a confidence-involved LiDAR association. Being economical and an end-to-end pipeline, multi-camera trackers [17, 18] receive attention and applications. Following the Tracking-By-Attention (TBA) framework, these trackers implement spatial-temporal modeling based on self/cross-attention and implicitly learn trajectory motion in a self-refinement paradigm. The trade-off between computational overhead, real-time performance, and tracking accuracy is the primary research direction in the current landscape.
@@ II-B Tracking-By-Detection
Estimation. This module utilizes filters (Linear Kalman Filter [3, 2, 4, 12, 11, 10], Extended Kalman Filter [1], Point Filter [14, 19], etc.), along with motion models, to perform two functions: (1) Predict the alive tracklets to achieve temporal alignment and association with the detection. (2) Update the matched tracklets with the corresponding observations, preparing prior information for downstream. Nevertheless, the real-time challenge emerges when managing numerous tracklets, as the filter involves heavy matrix operations on each tracklet. Therefore, we decouple and filter time-invariant states using a lightweight filter to alleviate the computational burden.
Association. As the core of the system, this module establishes tracklet-observation similarity and resolves matching correspondence. Geometry-based (IoU [10, 11], GIoU [2, 1, 4], Euclidean [14, 12, 3], NN distance [15, 6, 7, 5], etc.) and appearance-based are the commonly used affinity metrics. The former modeling process, reliant solely on spatial information, is mostly unsupervised and robust to occlusion. The latter, robust to inaccurate depth, can describe distant objects. Under the one-to-one assumption, the optimal assignment [28] algorithm is then used to obtain matching pairs. However, rotation in 3D space impairs the parallelization of IoU-like metrics, introducing severe latency. Fast-Poly proposes to leverage alignment to solve this dilemma.
Lifecycle. This module initializes, terminates, and merges tracklets based on the count-based strategy [3, 2, 1, 4, 11, 10] or the confidence-based [14, 1, 4, 12] strategy. The former has simple logic and strong robustness. The latter is more flexible in processing based on heuristic functions. In response to the inherent rigidity issue of the former approach, we employ score refinement to flexible terminate trajectories. In addition, we leverage trajectory average scores and a novel prediction to bolster occlusion robustness and enhance the estimation accuracy of the confidence-based strategy.
Figure 2: The pipeline of our proposed method. Our structure design is illustrated in Section III-A. Real-time improvements to the baseline [1] are highlighted in distinct colors. Orange denotes the Alignment to reduce the computational complexity of affinity calculations for rotated objects. Blue denotes the Densification to increase computational efficiency. Cyan denotes the Parallelization to execute pre-processing and motion prediction simultaneously, enhancing computational efficiency.
@@ III-A Overall Architecture
Despite sharing a similar pipeline, Fig. 2 illustrates the distinctions between our proposed method and Poly-MOT [1]. At each frame $t$, Fast-Poly employs two independent computing processes (Section III-D) to filter the 3D detections $D_{t}$ and predict existing trajectories $T_{t,t-1}$. Specifically, SF and NMS filters are leveraged to process the detections. Besides, the motion (time-variable), score (Section III-C1), and time-invariant (Section III-C3) states of trajectories are predicted by corresponding filters. $D_{t}$ and $T_{t,t-1}$ are then subjected to construct cost matrices for two-stage association. Notably, during NMS and matching, our proposed voxel mask (Section III-C2) and a novel geometry-based metric (Section III-B) accelerate computing. Hungarian algorithm [28] is then employed to obtain matched pairs $DT_{t}$, unmatched detections $D_{t}^{um}$, unmatched tracklets $T_{t-1}^{um}$. To relieve the matrix dimension, time-invariant and time-variable states in matched tracklets $T_{t-1}^{m}$ are updated with corresponding observations $D_{t}^{m}$ based on our proposed lightweight filter (Section III-C3) and Extended Kalman Filter (EKF), respectively. $T_{t-1}^{m}$ scores are also refined leveraging the confidence-count mixed lifecycle (Section III-C1). $D_{t}^{um}$ are then initialized as newly active tracklets. To flexibly determine FP agents in $T_{t-1}^{um}$, we soft-terminate the mismatch tracklets by considering their max-age and online average refined scores. The remaining tracklets $T_{t}$ are sent to downstream tasks and implemented for next frame tracking.
Figure 3: Top: Affinity difference between $A\text{-}gIoU$ and $gIoU$ for identical bbox pairs under varying geometric conditions in the 3D space. (a): overall affinity difference. (b): $IoU$ difference. (c): spatial occupancy difference. Bottom-Left: The calculation process of distinct metrics in the BEV space. Bottom-Right: The contributions of CGR across all trajectories within randomly sampled scenes from nuScenes of $A\text{-}gIoU$ and $gIoU$, employing our tracker with CenterPoint [19] detector.
@@ III-B Alignment
This section details how Fast-Poly mitigates high latency caused by object rotations through alignment, covering implementation and principles.
@@
A-GIoU. Varied from the pixel plane, the rotation causes overlap and convex hull irregular polygons instead of axis-aligned rectangles. The complexity of the well-known Sutherland-Hodgman [29] and Graham scan [30] algorithm are $\mathcal{O}(n^{2})$ and $\mathcal{O}(n\log{n})$, where $n$ is the number of input points. In addition, the sum calculation time increases rapidly with the growth of detections and tracklets. On the contrary, the solution complexity of the axis-aligned bounding box is only $\mathcal{O}(n)$. To this end, we proposed the Aligned Generalized Intersection over Union ($A\text{-}gIoU$), which is formulated as:
@@
where $B^{A}$ represents the axis-aligned $B$, containing the top-left and bottom-right points $p\in\mathbb{R}^{1\times 4}$ on the ground plane. Using the Bird’s-Eye View (BEV) space as a case study, the alignment procedure is depicted in Fig. 3(d). The overlap and convex hull calculations in Eq. 2 are consistent with $gIoU_{2d}$ [31], with $\mathcal{O}(n)$ computational complexity. For clarity in subsequent discussions, we designate the parts of $gIoU$ and $A\text{-}gIoU$ excluding $IoU$ as the spatial occupancy difference.
Delving into A-GIoU. As shown in Table II and V, on nuScenes, integrating $A\text{-}gIoU$ reduces latency by 73% while enhancing accuracy (+0.5% MOTA). This prompts the question: why does it excel in tracking tasks? A critical insight is that $A\text{-}gIoU$ aligns closely with $gIoU$ in describing the similarity between identical object pairs and quantifying the discrepancy between different matching instances.
Take 3D space as an example, firstly, Fig. 3(a) provides intuitive evidence of the similarity consistency between these two metrics for the same bbox pair across various geometric scenarios  [^2] Both bboxes use nuScenes Car average 3D size, with Z-position fixed at 1m and yaw angle sampled at $1.2^{\circ}$ step in $\in[0^{\circ},180^{\circ}]$. X-Y positions: Bbox1 at (0m, 0m); Bbox2 is sampled at 1m step in $[-5\text{m},5\text{m}]$.. Generally, $A\text{-}gIoU$ slightly exceeds $gIoU$, with tiny differences in small object overlaps (distant or parallel bbox pairs) cases. To elucidate these phenomena, Fig. 3(b, c) breaks down the distinctions between the two components of both metrics. Fig. 3(c) reveals that spatial occupancy difference is the key factor of this consistency. Fig. 3(b) further highlights the source of the minor distinction and the enhancement provided by $A\text{-}gIoU$: It increases the $IoU$ for closely positioned bbox pairs, improving the tracker to recall observations with strong positional correlation despite inaccurate heading angles.
Secondly, to validate the consistency between $A\text{-}gIoU$ and $gIoU$ in distinguishing true matches from false matches in data association, we introduce the Cost Gap Ratio (CGR) to quantify the performance disparity. CGR is defined as:
$CGR=(\min(C_{um})-C_{m})/(\max(C_{um})-C_{m}),$ (3)
where $C_{m}$ and $C_{um}$ are the cost of the true match and the false matches for a single tracklet  [^3] Outliers ($C_{m}\geq\min(C_{um}$)) are excluded (representing $\leq$ 3% of data) to constrain CGR within the (0, 1] interval.. A higher CGR indicates a more pronounced difference between $C_{m}$ and $C_{um}$, suggesting a higher probability of a true positive match. Fig. 3(e) illustrates CGR values across all trajectories within randomly sampled scenes from nuScenes. The consistency between $A\text{-}gIoU$ and $gIoU$ in characterizing distinct matching instances is confirmed by similar Quartile1 (Q1), median, and Quartile3 (Q3) values. In addition, for $A\text{-}gIoU$, at least 75% of CGR (Q1) exceeds 0.4 in the sampled scenes, ensuring the matching algorithm constructs accurate and stable bipartite graphs.
@@ III-C Densification
In this section, we improve the computational efficiency of the 3D tracking system through three key enhancements: soft tracklet elimination using tracking scores, distant similarity calculation prevention via the voxel mask, and time-invariant states splitting with a lightweight filter.
Confidence-Count Mixed Lifecycle. Different from the rigid The classic count-based lifecycle seriously raises tracker maintenance costs. In contrast, [14] flexibly terminates tracklets by manipulating scores, significantly reducing FP. However, as evidenced in [14] and Table VI, the deletion based on the latest score and the application of minus prediction detrimentally affect mismatch tracklets due to occlusion or FN detections. Towards this, a power function is first implemented to predict the scores smoothly. The refinement is described in two steps as follows:
@@
where $s_{t-1}$, $s_{t,t-1}$ and $s_{t}$ are the $t-1$ frame posterior score, $t$ frame priori score and $t$ frame posterior score of current tracklet, respectively. $c_{t}$ is the corresponding detection score. $\sigma$ is the hand-crafted decay rate, empirically $\leq 0.7$. The update function is similar with [14], satisfies that $s_{t}$ is more confidence than $s_{t,t-1}$ or $c_{t}$. After refinement, tracklets are terminated if their online average scores (computed from the initial frame to the current frame) fall below the deletion threshold $\theta_{dl}$ or if the number of mismatch frames exceeds the max-age. There are two noteworthy benefits: (1) Tracklets demonstrate increased resilience to temporary occlusions since their average confidence score decays gradually, as depicted in Table VI. (2) The power function swiftly pulls the tracklet predict score to the $\left[0,\sigma\right]$ interval, leaving the over-confidence area ($\left[0.7,1\right]$ interval in Fig. 2 of [14]). This reduces the hysteresis in score estimates, improving the tracking performance.
Voxel Mask. In 3D MOT, an overlooked consideration is: redundancy and matching are improbable between objects separated by significant distances in 3D Euclidean space. This starkly contrasts inherently no-depth 2D MOT [13, 20]. To address this issue, we propose the voxel mask inspired by the coarse-to-fine idea [32]. Specifically, before computing the global cost during data association and NMS iteration, we employ low-overhead Euclidean distance (2-norm) $||\cdot||_{2}$ and a hand-crafted size parameter $\theta_{vm}$ to generate voxel mask. The process can be expressed as:
@@
where $\mathcal{B}$ is the 3D bounding boxes representing detections or tracklets based on module representation. $\mathcal{B}^{xyz}$ are the 3D positions. $M_{t}$ is the voxel mask between $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. $\text{Aff}(\cdot)$ is the affinity function, constructing inter-object similarities between objects in pre-processing and association to detect redundancy and recall observations, respectively. Notably, $\text{Aff}(\cdot)$ directly pads invalid values into false index within $M_{t}$ to expedite, ultimately yielding $C_{t}$.
Lightweight Filter. Current 3D MOT methods embed the time-invariant states (Z-position $z$, width $w$, length $l$, height $h$) of trajectory into the Kalman Filter. However, the time invariance makes the Jacobian matrices of process and observation identity matrices, resulting in numerous meaningless 0-related calculations. Besides, the computational time escalates swiftly with an expanding number of trajectories. To this end, we decouple these states $X^{ti}$ from the KF and estimate them using a compact filter. The estimate process is expressed as:
@@
where $X^{ti}_{t-1}$, $X^{ti}_{t,t-1}$ and $X^{ti}_{t}$ are the $t-1$ frame posterior $X^{ti}$, $t$ frame priori $X^{ti}$ and $t$ frame posterior $X^{ti}$ of current tracklet, respectively. $D^{i}_{t}$ is the corresponding detection. LW-Filter($\cdot$) is the custom lightweight filter, implemented by a median or mean filter with length $l_{lw}$ in Fast-Poly. It stores nearest fixed-size time-invariant trajectory observations for prompt state estimation. There are two key insights: (1) LW-Filter($\cdot$) bypasses invalid computations, enabling fast filtering while preserving temporal data. (2) In contrast to KF, which entails manual process and observation noise, LW-Filter($\cdot$) relies solely on a single hyperparameter $l_{lw}$, reducing overfitting and ensuring robustness, as proven in Fig. 5.